OpenStudy (autumn9913):

Two groups of 10 soybean plants are measured for height in meters. Group A receives an experimental fertilizer while Group B receives the normal fertilizer treatment. Group A: 1.4, 1.3, 1.5, 1.5, 1.7, 1.3, 1.4, 1.3, 1.5, 1.4 Group B: 1.2, 1.3, 1.0, 1.2, 1.3, 1.4, 1.4, 1.2, 1.3, 1.3 Find the mean of Group A and B and the mean difference. Group A mean: _____ Group B mean: ______ Difference of means: ______ The groups are re-randomized and the mean differences are represented in the dot plot. Is the mean difference between the experimental and normal fertilizer significant?

OpenStudy (francesca_marie):

Is there a chart/graph with this?

OpenStudy (autumn9913):

here is the chart

OpenStudy (francesca_marie):

Oh okay hold on

OpenStudy (francesca_marie):

Group A mean = (1.4 + 1.3 + 1.5 + 1.5 + 1.7 + 1.3 + 1.4 + 1.3 + 1.5 + 1.4) / 10 = 1.43 Group B mean = (1.2 + 1.3 + 1.0 + 1.2 + 1.3 + 1.4 + 1.4 + 1.2 + 1.3 + 1.3) / 10 = 1.26 Difference of means = 1.43 - 1.26 = 0.17

OpenStudy (autumn9913):

Thanks